# Sorting inversion sequences

### v1

@@ Abstract
We consider the avoidance of patterns in inversion sequences that relate sorting via sorting machines including data structures such as pop stacks and stacks. Such machines have been studied under a variety of additional constraints and generalizations, some of which we apply here. We give the classification of several classes of sortable inversion sequences in terms of pattern avoidance. We are able to provide an exact enumeration of some of the sortable classes in question using both classical approaches and a more recent strategy utilizing generating trees.
@@ 1.   Introduction
An inversion in a permutation $\pi=\pi_{1}\pi_{2}\pi_{3}\cdots\pi_{n}$ is a pair of entries $\pi_{i},\pi_{j}$ where $i<j$, but $\pi_{i}>\pi_{j}$. An inversion sequence of length $n$ is a word $e=e_{0}\dotsb e_{n}$ which satisfies for each $i\in[n]=\{0,1,\ldots,n\}$ the inequalities $0\leq e_{i}\leq i$. These sequences correspond to permutations via a variant of the Lehmer code [17, 18] by mapping each entry $\pi_{j}$ of the permutation $\pi$ to $e_{j}$ where $e_{j}$ is the number of inversions $\pi_{j}$ is a part in $\pi$ as the second, smaller entry of the pair. The set of inversion sequences of length $n$ is denoted ${\mathbf{I}}_{n}$.
The study of pattern avoidance in inversion sequences began with work by Corteel, Martinez, Savage, and Weselcouch [8] and Shattuck and the first author [21]. Since then, extensive work has been done in this area, including recent work by Kotsireas, Yıldırım, and the first author [16] that introduces an algorithmic technique involving generating trees to determine enumeration of many pattern classes of inversion sequences (for example, see [6, 19, 20, 30] and the references therein). We utilize this algorithm ourselves in Section 4 to obtain a generating function for the number of inversion sequences sortable by two particular machines.
However, there remain many open problems when it comes to characterizing and enumerating the inversion sequences that avoid a class of consisting of a small number of relatively short (length less than five) patterns. Our interest is in the characterization and enumeration of inversion sequences sortable by stacks and pop stacks. We also study the characterization of inversion sequences sortable by data structures obtained by generalizing or further restricting the pop stacks.
A stack is a last-in, first-out sorting device with push and pop operations. Knuth [15] showed the permutation $\pi$ can be sorted (that is, by applying push and pop operations to the sequence $\pi_{1},\dots,\pi_{n}$ one can output the identity $1,\dots,n$) exactly if $\pi$ avoids the permutation $231$.
@@
Describing the permutations that are sortable by these machines is a point of interest in the field. To this end, it often possible to describe the sortable machine in terms of its basis.
@@
A restricted version of a stack that has received a fair amount of attention is a pop stack, first introduced by Avis and Newborn [4]. A pop stack has the same push and pop operations as a stack, but whenever a pop operation occurs, every entry in the pop stack is popped immediately. One foundational result is that the permutations that can be sorted by a single pop stack are exactly those that avoid $231$ and $312$. Some later works involving pop stacks include [1, 3, 7, 9, 12, 14, 25, 28].
Atkinson [2] explored a variation of a stack called an $(r,s)$-stack which allowed for an extension of the traditional push and pop operations. Specifically, the first index of an $(r,s)$-stack allowed for one to push an entry not only to the top of the stack, but to any of the top $r$ spots. Similarly, the second index expanded the options of the pop operation to allow one to pop any of the top $s$ entries from the stack. Thus a $(1,1)$-stack is simply a traditional stack. Atkinson was able to characterize these permutations when one of $r,s$ was fixed at $1$ and when $r=s=2$. Natural variations on stacks arise when studying inversion sequences that are similar in flavor to these generalized stacks.
Murray [10] considered restricting the depth of a stack, that is the maximum number of entries allowed in a stack at any given time, in the context of generating permutations with two stacks in series. Murray later continued this study with Lee and Rechnitzer [13] and then extended to sorting and pop stack sorting with Goh [11, 12]. While this work focused on a permutation sorting (or generating) machine consisting of two stacks in series, we will consider a similar restriction for a single stack or pop stack when sorting inversion sequences.
@@ 2.   Stack sorting inversion sequences
We give the standard optimal stack sorting algorithm in Algorithm 1.
@@
$p=p_{1}p_{2}\cdots p_{n}$ is a permutation of length $[n]$.
@@
         else if $p_{i}<$ top element of $S$ then
              push $p_{i}$ to the top of $S$
@@
For a word $w$ to be stack sortable, $w$ must avoid $120$, the same as avoiding $231$ to be able to sort permutation. The repetition allowed in words does not change this requirement from being both necessary and sufficient and so the proof does not change either. It is included below for completeness.
@@
The first $2$ of a $120$ pattern forces the smaller entry $1$ to the output before the 00 even enters the stack which creates an inversion with the $1,0$. Hence words containing a $120$ pattern are not stack sortable.
Next consider a word $w$ that is not stack sortable. The output $S(w)$ must contain an inversion, say an entry $b>a$ that appears before $a$. For $a$ to not have exited the stack before $b$, we know $a$ appears after $b$ in $w$. Further, for $b$ to have been forced out of the stack before $a$ entered, there must have been a larger entry $c$ that appeared after $b$, but before $a$ to force $b$ out before $a$ could enter the stack. Thus $w$ contains a $120$ pattern $bca$. ∎
@@
The enumeration of these sortable inversion sequences has been studied, but thus far remains open. See [21, Section 4] and [8] for initial study and further work on avoiding $120$ in [30].
@@ 3.   Pop stack sorting inversion sequences
The optimal pop stack sorting algorithm is mostly the same procedure as that for stacks, but when a pop stack is popped, all entries must exit the stack as described in Algorithm 2.
@@
$p=p_{1}p_{2}\cdots p_{n}$ is a permutation of length $[n]$.
@@
         else if $p_{i}<$ top element of $PS$ then
              push $p_{i}$ to the top of $PS$
@@
One can check that each of $120,201,$ and $1010$ are not pop stack sortable. As such, a word that contains any of these patterns cannot be sorted by a pop stack.
Conversely, suppose $w$ is not pop stack sortable. Consider what happens to force an entry to the output that is larger than an entry that appears later in the permutation. Using Algorithm 2, we had an entry $d$ that appears (immediately) after a smaller entry $b$ which forces $b$ out before $d$ can enter. If $b$ is the larger entry forced out, then there exists a smaller entry $a$ after $d$ and $bda$ form a $120$ pattern. If not, then there exists an entry $c>b$ that entered the pop stack before $b$ and there is an entry $a$ that appears after $d$ such that $c>a\geq b$. Note if $a>b$, then $cba$ form a $201$ pattern. Otherwise, if $a=b$, possible patterns $cbda$ are $2010,10$ 2010 , 1020, but only $1010$ does not contain one of the smaller forbidden patterns $120,201$. ∎
@@
The inversion sequences that are pop stack sortable are exactly those that avoid $120,201,$ and $1010$.
The pop stack sortable permutations are also known as layered permutations which are permutations of the form $\pi=\sigma_{1}\sigma_{2}\cdots\sigma_{k}$ where each layer $\sigma_{i}$ is a maximal length contiguous decreasing sequence such that all of the entries of $\sigma_{i}$ are less than all of the entries of $\sigma_{i+1}$. For example, the permutation $\pi=543216987$ is a layered permutation with layers $\sigma_{1}=54321,\sigma_{2}=6,\sigma_{3}=987$ as shown in Figure 1.
Figure 1: The layered permutation $\pi=543216987$
@@
A layered word $w$ is such that $w=\tau_{1}\tau_{2}\cdots\tau_{k}$ where each $\tau_{i}$ is
@@
For example, the word $w=211033275$ is a layered word with layers $2110,332,75$ shown in Figure 2.
Figure 2: The layered word $\pi=211033275$
@@
Now suppose $w$ avoids $120,201,1010$. Then $w$ is pop stack sortable. Consider the output from sorting $w$ broken into segments based on the pops. Reversing those segments reverts the word back to $w$. Further, by the algorithm, these breaks occur at exactly the strict ascents in $w$. The segments are then weakly decreasing and such that the smallest entry of any segment is at least as large as the largest entry of the previous segment for $w$ to be sortable and also maximal in length. Hence $w$ is layered. ∎
When restricting to inversion sequences, we are able to describe the generating tree which we describe in subsection 4.2. There are also some partial results we can obtain combinatorially based on the number of layers (which is also the number of pops needed to sort the sequence). Recall the Eulerian numbers $E(n,k)$ count the number of permutations of length $n$ with $k-1$ descents.
@@
In fact, the pop sortable inversion sequences of length $n$ with $k=1,2$ layers correspond exactly to the permutations of length $n$ with $k$ descents by the variant of Lehmer code used to create inversion sequences. In the trivial case of $k=1$, a permutation of length $n$ with $k-1=0$ descents is the identity permutation which is the only permutation of length $n$ to have an inversion sequence with one layer, namely the inversion sequence consisting of all 00s.
@@
As not all inversion sequences are pop stack sortable, this correspondence can clearly not continue indefinitely. Indeed, a permutation with even $k-1=2$ descents is not guaranteed a corresponding pop stack sortable inversion sequence. For example, the permutation $\pi=3214$ has the inversion sequence $e(\pi)=0120$ which is not sortable by a pop stack.
@@
A pop stack sortable inversion sequence of length $n\geq 1$ with $n-1$ layers will have exactly one entry $e_{i}$ as the second entry of a layer where $i\in\{2,3,4,\ldots,n\}$. It must be the case that $e_{m}=m-1$ for all $m<i$ for those entries to have all started new layers. This means $e_{i}=e_{i-1}=i-2$ or $e_{i}=e_{i-2}=i-3$. Further, we know $e_{i},e_{i+1},e_{i+2},\ldots,e_{n}$ are in increasing order as they are all in different layers.
@@
Otherwise, $e_{i}=i-3$ and so $i\geq 3$. Now because $i-2\leq e_{i+1}<e_{i+2}<\cdots<e_{n}\leq n-1$, up to two of these later entries could be two greater than the previous entry or one of these later entries could be three greater than the previous entry. For the cases when no entry is three more than the previous entry, we have $\binom{n-2}{1}+\binom{n-2}{2}+\binom{n-2}{3}$ ways to select the second entry of the length two layer and the entries that are two greater than the entry they follow. In the case where we have one entry that is three more than the previous entry, there are $\binom{n-2}{2}$ ways to select this entry and the entry that was the second in the longest layer. This gives us a total of $\binom{n-2}{1}+2\binom{n-2}{2}+\binom{n-2}{3}=\binom{n}{3}$ via a generalization of Pascal’s Identity.
@@
The number of pop sortable inversion sequences of length $n$ only containing $0,1$ is $(n-1)$st Cake number: $\binom{n-1}{0}+\binom{n-1}{1}+\binom{n-1}{2}+\binom{n-1}{3}$.

### v5

@@ Abstract
We consider the avoidance of patterns in inversion sequences that relate sorting via sorting machines including data structures such as stacks and pop stacks. Such machines have been studied under a variety of additional constraints and generalizations, some of which we apply here. We give the classification of several classes of sortable inversion sequences in terms of pattern avoidance. We are able to provide an exact enumeration of some of the sortable classes in question using both classical approaches and a more recent strategy utilizing generating trees.
@@ 1.   Introduction
An inversion in a permutation $\pi=\pi_{1}\pi_{2}\pi_{3}\cdots\pi_{n}$ is a pair of entries $\pi_{i},\pi_{j}$ where $i<j$, but $\pi_{i}>\pi_{j}$. An inversion sequence of length $n$ is a word $e=e_{1}\dotsb e_{n}$ which satisfies for each $i\in\{1,\ldots,n\}$ the inequalities $10\leq e_{i}\leq i-$ 1. These sequences correspond to permutations via a variant of the Lehmer code [21, 22] by mapping each entry $\pi_{j}$ of the permutation $\pi$ to $e_{j-1}$ where $e_{j-1}$ is the number of inversions in $\pi$ where $\pi_{j}$ is the second, smaller entry of the pair. The set of inversion sequences (respectively the set of inversion sequences of length $n$) is denoted ${\mathbf{I}}$ (respectively ${\mathbf{I}}_{n}$).
Example 1.1.
The permutation $\pi=51743862$ corresponds to the inversion sequence $e=01023026$.
The study of pattern avoidance in inversion sequences began with the work of Corteel, Martinez, Savage, and Weselcouch [12] as well as that of Shattuck and the first author [25]. Since then, extensive work has been done in this area, including [23, 24, 37] to list a few. Recently, Kotsireas, Yıldırım, and the first author [20] introduced an algorithmic technique involving generating trees to enumerate many pattern classes of inversion sequences (also see [7, 8] and the references therein). We utilize this algorithm in Section 4 to obtain a generating function for the number of inversion sequences sortable by two particular machines and give the succession rules to build the generating tree for a third machine.
There remain many open problems in characterizing and enumerating the inversion sequences that avoid a class consisting of a small number of relatively short (length less than five) patterns. Our interest is in the characterization and enumeration of inversion sequences sortable by stacks and pop stacks. We also study the characterization of inversion sequences sortable by data structures obtained by generalizing or further restricting the pop stacks.
A stack is a last-in, first-out sorting device with push and pop operations. Knuth [19] showed the permutation $\pi$ can be sorted by a stack (that is, by applying push and pop operations to the sequence $\pi_{1},\dots,\pi_{n}$ one can output the identity permutation $1,\dots,n$) if and only if $\pi$ avoids the permutation $231$.
@@
Describing the permutations that are sortable by machines consisting of one or more data structures (such as stacks or pop stacks) is a point of interest in the field. To this end, it is often possible to describe the sortable machine in terms of its basis.
@@
In the same vein, for any set of words $R$, denote by ${\mathbf{I}}(R)$ the set of inversion sequences in ${\mathbf{I}}$ that avoid every word in $R$. Similarly, denote by ${\mathbf{I}}_{n}(R)$ the set of inversion sequences of length $n$ that avoid every word in $R$.
@@
A restricted version of a stack that has received a fair amount of attention is a pop stack, first introduced by Avis and Newborn [5]. A pop stack supports the same push and pop operations as a standard stack, but whenever a pop operation occurs, every entry in the pop stack is popped immediately. One foundational result is that the permutations that can be sorted by a single pop stack are exactly those that avoid $231$ and $312$. Some later works involving pop stacks include [1, 4, 9, 11, 13, 16, 18, 30, 34].
Atkinson [2] explored a variation of a stack called an $(r,s)$-stack which extends the traditional push and pop operations. Specifically, the first index of an $(r,s)$ stack allowed one to push an entry not only to the top of the stack but to any of the top $r$ spots. Similarly, the second index expanded the options of the pop operation to allow one to pop any of the top $s$ entries from the stack. Thus a $(1,1)$-stack is simply a traditional stack. Atkinson was able to characterize these permutations when one of $r,s$ was fixed at $1$, as well as in the case when $r=s=2$. Natural variations on stacks arise when studying inversion sequences that are similar in flavor to these generalized stacks.
Elder [14] considered restricting the depth of a stack, that is the maximum number of entries allowed in a stack at any given time, in the context of generating permutations with two stacks in series. Elder later continued this study with Lee and Rechnitzer [17] and then extended to sorting and pop stack sorting with Goh [15, 16]. Although this work focused on a permutation sorting (or generating) machine consisting of two stacks in series, we will consider a similar restriction for a single stack or pop stack when sorting inversion sequences.
Note that natural sorting machines consisting of stacks and/or pop stacks do not give any advantage in sorting a sequence by inserting extra entries that might force a stack or pop stack to be popped at an earlier stage. As such, the sortable words from any such machine form a class and can be completely characterized by classical pattern avoidance.
@@ 2.   Stack sorting inversion sequences
We give the standard optimal stack sorting algorithm extended to words in Algorithm 1. Note that a word is considered sorted if all of its entries appear in weakly increasing order.
@@
$w=w_{1}w_{2}\cdots w_{n}$ is a word of length $[n]$.
@@
         else if $w_{i}\leq$ top element of $S$ then
              push $w_{i}$ to the top of $S$
@@
For a word $w$ to be stack sortable, $w$ must avoid $120$, the same as avoiding $231$ which is required for permutations to be sortable. The repetition allowed in words does not change this requirement from being both necessary and sufficient and so the proof does not change either. It is included below for completeness.
@@
The $2$ of a $120$ pattern forces the smaller entry $1$ to the output before the 00 even enters the stack which creates an inversion with the $1,0$. Hence words containing a $120$ pattern are not stack sortable.
Next consider a word $w$ that is not stack sortable. The output $S(w)$ must contain an inversion, say an entry $b>a$ that appears before $a$. For $a$ to not have exited the stack before $b$, we know $a$ appears after $b$ in $w$. Furthermore, for $b$ to have been forced out of the stack before $a$ entered, there must have been a larger entry $c$ that appeared after $b$, but before $a$ to force $b$ out before $a$ could enter the stack. Thus $w$ contains a $120$ pattern $bca$. ∎
@@
The enumeration of these sortable inversion sequences has been studied, but thus far remains open. See [25, Section 4] and [12] for initial study and further work on avoiding $120$ in [37].
@@ 3.   Pop stack sorting inversion sequences
The optimal pop stack sorting algorithm is mostly the same procedure as that for stacks, but when a pop stack is popped, all entries must exit the stack as described in Algorithm 2. Note that the end result of this algorithm is a reversal of each maximal weakly descending run of the input sequence.
@@
$w=w_{1}w_{2}\cdots w_{n}$ is a word of length $[n]$.
@@
         else if $w_{i}\leq$ top element of $PS$ then
              push $w_{i}$ to the top of $PS$
@@
We note that Cerbai [9, Theorem 5.2] found the basis for sortable Cayley words in the context of a hare pop stack which is simply a pop stack that only allows entries to appear in weakly decreasing order from top to bottom (which is a necessary requirement for sorting with a single stack). This basis is the same as the one in Theorem 3.1 below. Cerbai’s argument could also be extended to sorting general words by a pop stack as we have done here.
@@
One can check that none of $120,201,$ and $1010$ are pop stack sortable. As such, a word that contains any of these patterns cannot be sorted by a pop stack.
Conversely, suppose $w$ is not pop stack sortable. Then the image of $w$ under the pop stack operator, which we will denote by $PS(w)$, must contain an inversion $(y,x)$. That is, $y>x$ and $y$ appears before $x$ in $PS(w)$. Consider what happens to force the entry $y$ to the output before $x$.
Using Algorithm 2, there must be a point when $y$ is in the stack and an entry $d$ that appears (immediately) after a smaller entry $b$ in $w$. Because $b<d$ and $b$ appears on the top of the pop stack when $d$ is the next entry of $w$, the pop stack must be popped before $d$ can enter. As $x$ appears after $y$ in $PS(w)$, it must be that $x$ is still in the input at this state. These entries (if they are all distinct) are shown before entering the pop stack (left) and in the state where $b$ just entered the pop stack (right) in Figure 1.
If $y=b$, then $x$ appears after $d$ in $w$. Thus $bdx$ forms a $120$ pattern in $w$.
Otherwise, there exists an entry $y>b$ that is also forced out before $d$ enters the pop stack but before $x$. Thus, $y$ appears before $b$ in $w$. Either we can assume $x=d$ (that is, $y>d$) or if $x<d$, then $x$ appears after $d$ in $w$. In the former case, $ybd$ is a $201$ pattern in $w$. In the latter case, $x<d$ appears after $d$, so consider the pattern $ybdx$ in $w$ where we are reduced to the options $d\geq y>x,b$. Thus, $ybdx$ is one of the following patterns $1010,1020,2021,2120,2031,2130$. Of these patterns, only $1010$ does not contain one of the smaller forbidden patterns $120,201$. ∎
[…]
Figure 1: Moving to the state where $y$ will be forced to exit the pop stack before $x$ enters
@@
The inversion sequences that are pop stack sortable are those in ${\mathbf{I}}(120,201,1010)$.
The pop stack sortable permutations are also known as layered permutations which are permutations of the form $\pi=\sigma_{1}\sigma_{2}\cdots\sigma_{k}$ where each layer $\sigma_{i}$ is a maximal length contiguous decreasing sequence such that all of the entries of $\sigma_{i}$ are less than all of the entries of $\sigma_{i+1}$. For example, the permutation $\pi=543216987$ is a layered permutation with layers $\sigma_{1}=54321,\sigma_{2}=6,\sigma_{3}=987$ as shown on the left in Figure 2.
$\pi=543216987$$w=211033275$
Figure 2: The layered permutation $\pi=543216987$ and the layered word $w=211033275$.
@@
A layered word $w$ is a word $w=\tau_{1}\tau_{2}\cdots\tau_{k}$ where each $\tau_{i}$ is
@@
For example, the word $w=211033275$ is a layered word with layers $2110,332,75$ shown on the right in Figure 2.
@@
Now suppose $w$ avoids $120,201,1010$. Then $w$ is pop stack sortable. Consider the output from sorting $w$ broken into segments based on the pops. Reversing those segments reverts the word back to $w$. In addition, according to the algorithm, these breaks occur exactly at the strict ascents in $w$. The segments then are weakly decreasing so that the smallest entry of any segment is at least as large as the largest entry of the previous segment for $w$ to be sortable and also maximal in length. Hence $w$ is layered. ∎
In Section 4.2, we are able to describe the generating tree for layered inversion sequences. There are also some partial results we can obtain combinatorially based on the number of layers (which is also the number of pops needed to sort the sequence). Recall that the Eulerian numbers $E(n,k)$ count the number of permutations of length $n$ with $k-1$ descents.
@@
In fact, the pop sortable inversion sequences of length $n$ with $k=1,2$ layers correspond exactly to the permutations of length $n$ with $k-1$ descents by the variant of Lehmer code used to create inversion sequences. In the trivial case of $k=1$, a permutation of length $n$ with $k-1=0$ descents is the identity permutation which is the only permutation of length $n$ to have an inversion sequence with one layer, namely the inversion sequence consisting of all 00s.
@@
As not all inversion sequences are pop stack sortable, this correspondence can clearly not continue indefinitely. Indeed, a permutation with even $k-1=2$ descents is not guaranteed a corresponding pop stack sortable inversion sequence. For example, the permutation $\pi=3214$ has the inversion sequence $e(\pi)=0120$, which cannot be sorted by a pop stack.
@@
A pop stack sortable inversion sequence $e$ of length $n\geq 1$ with $n-1$ layers will have exactly one entry $e_{i}$ as the second entry of a layer where $i\in\{2,3,4,\ldots,n\}$. It must be the case that $e_{m}=m-1$ for all $m<i$ for those entries to have all started new layers. Note that $e_{i}\geq e_{i-2}$ as otherwise $e_{i-2},e_{i-1},e_{i}$ would form a forbidden $120$ pattern in $e$. This means $e_{i}=e_{i-1}=i-2$ or $e_{i}=e_{i-2}=i-3$. Further, we know $e_{i},e_{i+1},e_{i+2},\ldots,e_{n}$ are in strictly increasing order as they are all in different layers.
@@
Otherwise, $e_{i}=i-3$ and so $i\geq 3$. Now because $i-2\leq e_{i+1}<e_{i+2}<\cdots<e_{n}\leq n-1$, up to two of these later entries could be two greater than the previous entry or one of these later entries could be three greater than the previous entry. For the cases where no entry is three more than the previous entry, we have $\binom{n-2}{1}+\binom{n-2}{2}+\binom{n-2}{3}$ ways to select the second entry of the length two layer and the entries that are two greater than the entry they follow. In the case where we have one entry that is three more than the previous entry, there are $\binom{n-2}{2}$ ways to select this entry and the entry that was the second in the longest layer. This gives us a total of $\binom{n-2}{1}+2\binom{n-2}{2}+\binom{n-2}{3}=\binom{n}{3}$ via a generalization of Pascal’s Identity.
@@
The number of pop stack sortable inversion sequences of length $n$ only containing $0,1$ is the $(n-1)$st Cake number: $\binom{n-1}{0}+\binom{n-1}{1}+\binom{n-1}{2}+\binom{n-1}{3}$.
We note that Cerbai [9] also considered a tortoise pop stack for sorting Cayley words that had an additional pattern restriction of not allowing a $00$ pattern to appear in the pop stack. (Restricting the content of stacks in terms of pattern avoidance was initially considered in [3, 33, 10].) Cerbai’s characterization of sortable Cayley words can be easily extended to all words, so we omit the proof here.
Proposition 3.8.
The words (and thus inversion sequences) that are tortoise pop stack sortable are those which avoid $120,201,110,$ and $100$.
Given this classification, the generating function for the enumeration of tortoise pop stack sortable inversion sequences is given in Theorem 12 in recent work by Callan and Mansour [8].